OpenStudy (zeesbrat3):

The half-life of bismuth-210, 210Bi, is 5 days. (a) If a sample has a mass of 344 mg, find the amount remaining after 15 days. (b) Find the amount remaining after t days. I have already found the answer to part a but I cannot get the second part.

OpenStudy (firedragonrage):

whats the answer to the first part? then maybe I can do it.

OpenStudy (firedragonrage):

What you have to do is make an equation for the answer to A. That is then the answer to B.

OpenStudy (unklerhaukus):

The initial amount is \(A_0\), (the half-life is \(t_{1/2}\)) and the amount remaining after a time \(t\), is \(A(t)\) \[\large A(t) = A_0(\tfrac12)^{t/t_{1/2}}\]

OpenStudy (eliesaab):

The formula first 5 years, we get 344/2=172 the next 5 years we 172/2=86 the last 5 years we get 86/2 =43